A toroid has an inside diameter of 40.0 cm and an outside diameter of 45.0 cm. It conducts a current of 30.0 A in its 700 turns. The magnetic field at a distance of 42 cm from its center is:

a. 50 mT
b.30 mT
c.20 mT
d.10 mT
Answer #1

Solution)

Given,

Inside Diameter, D=40 cm

Radius, r=20*10^-3 m

Outside diameter=45 cm

Current, I=30 A

No of turns, N=700

Distance, r=0.42 m

We know,

Magnetic Field, B=uoNI/2pi*r

On substituting values,

B=((2*10^-7)*700*30)/0.42=0.1 mT (option d)------Ans

============